OpenStudy (anonymous):

How can x - 5 = x + 2 be set up as a system of equations?

OpenStudy (btaylor):

If you are looking for the solutions for 2 lines, write each side as a linear equation: y = x-5 y = x+2 DONE.

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

No wait... I just noticed the question was written wrong.

OpenStudy (anonymous):

1/6x - 5 = 1/5x + 2 Do I still set it up that same way?

OpenStudy (anonymous):

yup.

OpenStudy (anonymous):

okay thanks guys :D